### v1

@@ 3. Robustly Periodic chain-recurrence classes
Next, we present the aforementioned result.
Lemma 3.5.
If $X\in\mathcal{R}$ and $\Lambda\subset M$ is a non-trivial sectional-hyperbolic chain-recurrent class, then it is a homoclinic class.
Proof.
$\Lambda^{\prime}=\Lambda\cup\bigcup_{n\geq 1}\gamma_{n}.$
Claim 1.
$\Lambda^{\prime}$ is a compact and invariant subset of $U$.
Proof.
First, observe that $\Lambda^{\prime}\subset U$ and $\Lambda^{\prime}$ is invariant by definition. Next, let $\{x_{k}\}_{k\geq 1}\subset\Lambda^{\prime}$ be a sequence such that $x_{k}\to x\in M$, when $k\to+\infty$. If $x_{k}\in\Lambda$ for sufficiently large $k$, then $x\in\Lambda$ since $\Lambda$ is compact. Similarly, if $x_{k}\subset\gamma_{n}$ for sufficiently large $k$ and some $n\geq 1$ fixed, then $x\in\gamma_{n}\subset\Lambda^{\prime}$. It remains to verify the case where there are sequences $k_{l},n_{l}\to\infty$, $n_{l}\neq n_{l^{\prime}}$, for $l\neq l^{\prime}$ and $x_{k_{l}}\subset\gamma_{n_{l}}$. In this case, since $d^{H}(\gamma_{n},\Lambda)\to 0,$ we have $d(x_{k_{l}},\Lambda)\to 0$. Hence, $x\in\Lambda\subset\Lambda^{\prime}$ and the claim holds. ∎
Notice that by Lemma 2.7 the set $\Lambda^{\prime}$ is sectional hyperbolic. Now, we consider the following cases:
Case 1: $\Lambda$ does not contain singularities
First, let $U$ be the neighborhood given by Lemma 2.7 and let $\Lambda^{\prime}$ be defined as above. Since $\Lambda^{\prime}\subset U$, Claim 1 and the hyperbolic lemma imply that $\Lambda^{\prime}$ is a compact invariant hyperbolic set. Let $T{\Lambda^{\prime}}=E^{\prime s}\oplus\langle X\rangle\oplus E^{\prime u}$ be the hyperbolic splitting of $\Lambda^{\prime}$. It is well known that dominated splittings vary continuously and hence $\dim(E^{\prime s})$ and $\dim(E^{\prime u})$ are constant along $\Lambda$, because $\Lambda$ is a chain transitive set. Moreover, up to possibly reducing $U$, we can assume $dim(E^{\prime s})$ and $\dim(E^{\prime u})$ are constant along $\Lambda^{\prime}$ because $d^{H}(\gamma_{n},\Lambda)\to 0$. Also, the continuity of the splitting and the compactness of $\Lambda^{\prime}$ implies that the angles between the fibers $E^{\prime s}$, $\langle X\rangle$ and $E^{\prime u}$ are uniformly limited away from zero.
[…]
for $n$ sufficiently large. Consequently, $\Lambda$ contains a periodic orbit.
Case 2: $\Lambda$ contains singularities
Now we assume that $\Lambda=C(\sigma_{X})$, for some singularity $\sigma\in\Lambda$. We shall proceed similarly to Case 1, but here the proof is more delicate, since $C(\sigma_{X})$ is not hyperbolic. Again, consider the neighborhood $U$ from Lemma 2.7 and $\Lambda^{\prime}$. Thus, $\Lambda^{\prime}$ is sectional hyperbolic. Let $T\Lambda^{\prime}=E\oplus F$ be the sectional-hyperbolic splitting of $\Lambda^{\prime}$. Similarly to Case 1, the continuity of the splitting implies that $\dim(E)$ and $\dim(F)$ are constant along $\Lambda^{\prime}$ and that the angles between $E$ and $F$ are uniformly bounded away from zero. Since $E$ is uniformly contracted, by the classical stable manifold theory, for every $x\in\Lambda^{\prime}$ there is a well defined local stable manifold $W^{ss}_{loc}(x)$ with uniform size along $\Lambda^{\prime}$.
In contrast with Case 1, since $\Lambda^{\prime}$ is not hyperbolic, there may not be well defined for local unstable manifolds for every point in $\Lambda^{\prime}$. Moreover, even though these local unstable manifolds may exist for some points (e.g. along non-singular compact and invariant subsets) they do not need to have uniform size. To overcome this issue, we firstly recall that the periodic orbits $\gamma_{n}$ are hyperbolic by the hyperbolic lemma. So, fix $\gamma_{n}$ and let
$T\gamma_{n}=E^{\prime s}\oplus\langle X\rangle\oplus E^{\prime u}$
be the hyperbolic splitting of $\gamma_{n}$. Since $E$ is uniformly contracted, we have $E(x)\subset E^{\prime s}(x)$, for every $x\in\gamma_{n}$. Now, suppose there is $x\in\gamma_{n}$ so that $E^{\prime s}(x)\not\subset E(x)$. In this case, there is a vector $v\in E^{\prime s}(x)$ such that $v=v^{E}+v^{F}$, $v^{E}\in E(x)$, $v^{F}\in F(x)$ and $v^{F}\neq 0$. In particular $v^{F}\subset E(x)$. Recall that $\langle X(x)\rangle\subset F$ for any $x\in\gamma_{n}$. Since $\gamma_{n}\cap Sing(X)=\emptyset$, there are $A_{2}>A_{1}>0$ such that $A_{1}\leq X(x)\leq A_{2},$ for any $x\in\gamma_{n}$. But this implies that the flow direction along $\gamma_{n}$ can neither be uniformly contracted or expanded by $DX_{t}$. Consequently, $v^{F}\not\in\langle X(x)\rangle$. On the other hand, by considering the subspace $L\subset F$ generated by $v^{F}$ and $\langle X(x)\rangle$, the sectional-expansion of $L$ implies, $v^{F}\notin E^{\prime s}(x)$. This contradiction implies $E(x)=E^{\prime s}(x)$, for every $x\in\gamma_{n}$. Consequently,
$\langle X(x)\rangle\oplus E^{\prime u}(x)=F(x),$
for any $x\in\gamma_{n}.$
Let $r>0$ be such that $B_{r}(\sigma_{X})\cap Sing(X)=\{\sigma_{X}\}$. Since any singularity contained in $C(\sigma_{X})$ is hyperbolic by Lemma 2.2, there are only finitely many of them. By shrinking $U$ if it is necessary, we have that $Sing(X|_{U})=Sing(X|_{C(\sigma_{X})})$ and hence, there is a neighborhood $V$ of $Sing(X|_{C(\sigma_{X})})$ contained in $U$ satisfying
  • $V=V_{1}\cup\cdots\cup V_{j}$.
  • Each $V_{i}$ is a neighborhood of a unique singularity in $X|_{U}$.
  • $V_{i}\subset B_{r}(\sigma_{X})$, for any $i=1,\ldots,j$.
  • There is $0<L_{0}<1$ such that $\|X(z)\|<L_{0}$, for every $z\in V$.
Recall we are assuming $C(\sigma_{X})$ is non-trivial. Since $\gamma_{n}$ converges to $C(\sigma_{X})$ in the Hausdorff topology, then we can assume that each $\gamma_{n}$ contains a point $p_{n}$ so that $p_{n}\to\sigma_{X}$. So, by a classical Grobman-Hartman argument, there is a sequence of times $t_{n}^{+}>0$ so that, by denoting $p^{s}_{n}=\phi_{t_{n}^{+}}(p_{n})$, we can assume, by possibly taking subsequences, that
$p^{s}_{n}\to p^{u},$
where $p^{u}\in W^{u}(\sigma_{X})$. By possibly reducing $V$, we can assume $p^{u}\in B_{r}(\sigma_{X})\setminus V$. On the other hand, since $d^{H}(C(\sigma_{X}),\gamma_{n})\to 0$, there is a sequence of regular points $x^{u}_{n}\in C(\sigma_{X})$ such that $d(x^{u}_{n},p^{s}_{n})\to 0$. Then, $x^{u}_{n}\to p^{u}$. Hence, $p^{u}\in C(\sigma_{X})\setminus Sing(X|_{C(\sigma_{X})})$, because $C(\sigma_{X})$ is closed.
Let
$K_{0}=\inf\{|X(x)|:x\in U\setminus V\},$
and notice that $K_{0}>0$. Let $K_{1}>0$ be such that $K_{0}<|X(x)|<K_{1}$, for every $x\in U\setminus V$. Let $K,\lambda>0$ the constants given by sectional hyperbolicity of $\Lambda^{\prime}$, and let $T_{0}>0$ in a such way that
(2) $K_{2}e^{\lambda t}\geq e^{\lambda^{\prime}t}\text{ and }K_{3}^{2}e^{\lambda^{\prime}t}\geq e^{\lambda^{\prime\prime}t},\quad\forall t\geq T_{0},$
where
$K_{2}=\min\left\{K,\frac{K}{K_{1}}\right\}>0,\;K_{3}=\min\left\{\frac{K_{0}}{L_{0}},\frac{K_{0}}{K_{1}}\right\}>0\textrm{ and }0<\lambda^{\prime\prime}<\lambda^{\prime}<\lambda.$
Since $p_{n}\to\sigma$, one has $\tau_{n}\to\infty$, where $\tau_{n}$ is the period of $\gamma_{n}$. In particular, there is a natural number $k_{0}>0$ so that such that $\tau_{n}>T_{0}$, for every $n\geq N_{1}$.
Now, let $\eta=(e^{\lambda^{\prime\prime}T_{0}})^{-1}\in(0,1)$ and let $0=t_{0}<t_{1}<\cdots<t_{m}<\cdots$ such that $t_{i+1}-t_{i}=T_{0}$ for every $i=0,\ldots,k-1$. On one hand, notice that for every unitary vector $v_{i}\in\mathcal{F}_{X_{t_{i}}(p_{n}^{s})}$ one has
$Ke^{\lambda(t_{i+1}-t_{i})}$ $\leq$ $|\det\Phi_{t_{i+1}-t_{i}}(X_{t_{i}}(p_{n}^{s}))|$
$=$ $\frac{\|\Phi_{t_{i+1}-t_{i}}(X_{t_{i}}(p_{n}^{s}))\|\cdot\|X(X_{t_{i+1}}(p_{n}^{s}))\|\cdot\sin(\theta_{t})}{\|X(X_{t_{i}}(p_{n}^{s}))\|}$
$=$ $\frac{\|X(X_{t_{i+1}}(p_{n}^{s}))\|}{\|X(X_{t_{i}}(p_{n}^{s}))\|}\cdot\|\psi_{t_{i+1}-t_{i}}(X_{t_{i}}(p_{n}^{s}))v_{i}\|,$
where $\theta_{t}=\angle(\Phi_{t_{i+1}-t_{i}}(X_{t_{i}}(p_{n}^{s})),X(X_{t_{i+1}}(p_{n}^{s})))$, so that by the choice of $K_{2}$ in (2),
$\|\psi^{*}_{t_{i+1}-t_{i}}(X_{t_{i}}(p_{n}^{s}))|_{\mathcal{F}_{X_{t_{i}}(p_{n}^{s})}}\|$ $\geq$ $\|\psi^{*}_{t_{i+1}-t_{i}}(X_{t_{i}}(p_{n}^{s}))v_{i}\|$
$=$ $\left[\frac{\|X(X_{t_{i}}(p_{n}^{s}))\|}{\|X(X_{t_{i+1}}(p_{n}^{s}))\|}\right]^{2}|\det\Phi_{t_{i+1}-t_{i}}(X_{t_{i}}(p_{n}^{s}))|$
$\geq$ $\left[\frac{\|X(X_{t_{i}}(p_{n}^{s}))\|}{\|X(X_{t_{i+1}}(p_{n}^{s}))\|}\right]^{2}e^{\lambda^{\prime}(t_{i+1}-t_{i})},\quad\forall i=0,\ldots,k-1.$
Thus, since $X_{t_{k}}(p_{n}^{s})$ belongs either $V$ or $U\setminus V$, we have by the choice of $K_{3}$ in (2) that
$\prod_{i=0}^{k-1}\|\psi^{*}_{t_{i+1}-t_{i}}(X_{t_{i}}(p_{n}^{s}))|_{\mathcal{F}_{X_{t_{i}}(p_{n}^{s})}}\|$ $\geq$ $\prod_{i=0}^{k-1}\left[\frac{\|X(X_{t_{i}}(p_{n}^{s}))\|}{\|X(X_{t_{i+1}}(p_{n}^{s}))\|}\right]^{2}e^{\lambda^{\prime}(t_{i+1}-t_{i})}$
$\geq$ $\left[\frac{\|X(p_{n}^{s})\|}{\|X(X_{t_{k}}(p_{n}^{s}))\|}\right]^{2}e^{\lambda^{\prime}kT_{0}}$
$\geq$ $\left(K_{3}\right)^{2}e^{\lambda^{\prime}kT_{0}}$
$\geq$ $(e^{\lambda^{\prime\prime}T_{0}})^{k}=(\eta^{-1})^{k}.$
So, by Theorem 2.6, the local unstable manifold $W^{u}_{\delta_{1}}(p^{n}_{s})$ contains a segment of size $\delta_{1}\|X(p^{s}_{n})\|\geq\delta_{1}K_{0}$, with $\delta_{1}=\delta_{1}(\eta,T_{0})$, for any $n\geq N_{1}$. On the other hand, since $E$ is uniformly contracting on $\Lambda^{\prime}$ there are local stable manifolds $W_{\varepsilon_{0}}^{s}(p_{n}^{s})$ of uniform size and tangent to $E(p_{n})\oplus\langle X(p_{n})\rangle=E^{\prime s}(p_{n})\oplus\langle X(p_{n})\rangle$, we say $\varepsilon_{0}>0$, on $p_{n}^{s}$ for every $n\geq 1$. Hence, if we choose $\varepsilon=\min\{\varepsilon_{0},\delta_{1}\}$, we have by the convergence of $p^{s}_{n}$ to $p^{u}$ that there is a natural number $N>N_{1}$ such that
$W_{\varepsilon}^{u}(p^{s}_{n})\pitchfork W_{\varepsilon}^{s}(p_{m}^{s})\text{ and }W_{\varepsilon}^{s}(p^{s}_{n})\pitchfork W_{\varepsilon}^{u}(p_{m}^{s}),\quad n,m\geq N.$
We have proved that $p_{n}\in C(p_{N})$ for every $n\geq N$ and hence $p_{n}^{s}\in C(p_{N})$ for every $n\geq N$. To finish the proof we need to show than $\sigma_{X}\in C(p_{N})$. For this sake, recall that $p^{s}_{n}\to p^{u}\in C(\sigma_{X})\cap W^{u}(\sigma_{X})$. Since $W^{u}(\sigma_{X})$ is invariant and uniformly contracted in the past, we conclude $\sigma_{X}\sim p_{N}$. Analogously, since $p^{u}_{n}\to p^{s}\in C(\sigma_{X})\cap W^{s}(\sigma_{X})$ we conclude $p_{N}\sim\sigma_{X}$. This implies $C(\sigma)\supset C(p_{N})$ proving that $p_{N}\in C(\sigma)$. Then, the result is obtained by item (2) of Lemma 3.2. ∎
According to Lemma 3.5, it remains to prove that the continuation $\Lambda_{Y}$ of $\Lambda$ contains a periodic orbit. When $\Lambda$ contains singularities, the argument is divided into two steps:
@@
By Lemma 3.5, there are a neighborhood $V$ of $Sing(C(\sigma_{X}))$ a point $p^{s}\in C(\sigma_{X})\cap W^{s}(\sigma_{X})$ and a sequence of periodic orbits $\gamma_{n}\in C(\sigma_{X})$, $n\geq 1$, such that $p_{n}^{u}\in\gamma_{n}$ converges to $p^{s}$ when $n$ goes to infinity.
Let $\xi_{k}$, $k\geq 1$, be a decreasing sequence of positive numbers such that $\xi_{k}\to 0$. For every $k>0$, choose $N_{k}>0$ such that $d(p_{n}^{u},p^{s})<\xi_{k}$ for every $n\geq N_{k}$. Fix $\alpha>0$ as in Lemma 2.5, with respect to $r$, $C(\sigma)$ and $T_{0}$. Following the argument given in the proof of Case 2 of Lemma 3.5, there are $N\in\mathbb{N}$ and $\delta_{1}>0$ such that the local unstable manifold $W^{u}_{\delta_{1}}(p_{n}^{u})$ contains a segment of size $\delta_{1}\|X(p^{s}_{n})\|>\delta_{1}K_{0}$, with $\delta_{1}=(\eta,T_{0})$, for any $n\geq N$. Since $d(p_{n}^{u},p^{s})\leq\xi_{k}$, for every $n\geq N_{k}$ and $\xi_{k}\to 0$, we conclude there is a natural number $k_{1}\geq k_{0}$ such that $p^{n}_{u}\in B_{r}(\sigma)\cap\overline{D_{\alpha}^{cs}(\sigma)}$, for every $n\geq N_{k_{1}}$. So, Lemma 2.5 gives
(3) $W^{u}(p^{u}_{n})\pitchfork W^{s}_{loc}(\sigma_{X})\neq\emptyset,$
for every $n\geq N_{k_{1}}$.
@@
Remark 3.7.
From the above lemma we derive the following observations:
  • In the proof of Case 2, we connect a periodic orbit $\gamma$ of $X$ with the singularity $\sigma\in C(\sigma)$. Since the periodic orbits accumulate on $C(\sigma)$ in the Hausdorff topology, according to item 2 of Lemma 3.2, we deduce that (3) holds for $\gamma_{n}=\gamma$ and any singularity $\rho\in C(\sigma)$.
Now, recall that for a non singular point $x\in M$, the orthogonal complement $N_{x}$ of $X(x)$ is defined by
$N_{x}=\{v\in T_{x}M:\langle v,X(x)\rangle=0\}.$
Let $\Lambda=C(\sigma_{X})$ and $\Lambda_{Y}=C(\sigma_{Y})$. Fix a neighborhood $U$ for $\Lambda$ and $\Lambda_{Y}$, given by the continuity of $Y\in\mathcal{U}\to C(\gamma_{Y})$. Since $\Lambda$ is partially hyperbolic, it is possible to extend the SH splitting $T_{\Lambda}M$ to $U$ (by shrinking $U$ if it is possible), that we will keep written as $T_{U}M=E^{s}\oplus F$. It should be noted that since $\Lambda$ is sectional-hyperbolic one has $1\leq\dim E^{s}\leq n-2$ and $1\leq\dim F\leq n-1$. Thus, the subspace $N_{x}$ is of codimension one and has a coordinate system induced by the SH splitting for every $x\in U$ in the following way:
$\mathcal{E}^{s}_{x}=\pi(E_{x}^{s})\text{ and }\mathcal{F}_{x}=\pi(F)\text{ or }F_{x}\cap N_{x},\quad x\in U,$
where $\pi:TM\to N$ denotes the orthogonal projection. From partial hyperbolicity we have that the angle between these subspaces is uniformly bounded away from zero. In this way, denote by $N(\eta)$ the standard cube $(-\eta,\eta)^{n-1}\subset N_{x}$, according to this coordinate system.
For any point $x\in U$, there is a number $\eta(x)>0$ such that the exponential map $Exp_{x}:N_{x}\to M$ is an isometry. Note that for every $\delta>0$, there is $\eta_{\delta}>0$ such that $\eta(x)>\eta_{\delta}$ for any $x\in U\setminus B(Sing(\Lambda),\delta)$. Let $\eta^{*}:U\to\mathbb{R}^{+}$ such that $\eta^{*}(x)=\min\{\eta(x),\eta_{\delta}\}$ for every $x\in U$. Denote by $\mathcal{N}_{x}=Exp_{x}(N_{x}(\eta^{*}(x)))\subset M$, a codimension 1 cross section on $x$ of size $\eta^{*}(x)$.
For a regular point $x$, the coordinates $\mathcal{E}^{s}_{x}$ and $\mathcal{F}_{x}$ allows us to define boxes $B\subset N_{x}$ with respect to these axis as follows: Denote by $\hat{\pi}_{s}:N_{x}\to\mathcal{E}^{s}_{x}$ and $\hat{\pi}_{c}:N_{x}\to\mathcal{F}_{x}$ the projections along the $\mathcal{E}^{s}_{x}$-axis and the $\mathcal{F}_{x}$-axis respectively. The stable (central) volume of $B$ is given by
@@
where $vol_{s(c)}(A_{s(c)})$ denotes the volume of a subset $A\subset\mathbb{R}^{\dim\hat{E}_{x}^{s}(\hat{F}_{x})}$. We shall use a vector $v=(\varepsilon^{s}_{1},\varepsilon^{s}_{2},\ldots,\varepsilon^{s}_{n_{s}},\varepsilon^{c}_{1},\varepsilon^{s}_{2},\ldots,\varepsilon^{c}_{n_{c}})$, where $1\leq n_{s}\leq n-2$ and $1<n_{c}\leq n-1$ satisfy $n_{s}+n_{c}+1=n$, to describe the box $B[v]$ with stable volume $\varepsilon_{s}$ and central volume $\varepsilon_{c}$.
@@
$H_{x}^{t}:dom(G_{x}^{t})\subset\mathcal{N}_{x}\to\mathcal{N}_{X_{t}(x)},$

### v2

@@ 3. Robustly Periodic chain-recurrence classes
In order to prove Theorem 3.1 in the singular case, we divided the argument into two steps:
@@
$\Lambda^{\prime}=\Lambda\cup\bigcup_{n\geq 1}\gamma_{n}.$
Notice that by Claim 1 of Lemma 3.4 and Lemma 2.4 the set $\Lambda^{\prime}$ is sectional hyperbolic.
Now we assume that $\Lambda=C(\sigma_{X})$, for some singularity $\sigma_{X}\in\Lambda$. Consider the neighborhood $U$ from Lemma 2.4 and $\Lambda^{\prime}$. Let $T\Lambda^{\prime}=E\oplus F$ be the sectional-hyperbolic splitting of $\Lambda^{\prime}$. Note that the continuity of the splitting implies that $\dim(E)$ and $\dim(F)$ are constant along $\Lambda^{\prime}$ and that the angles between $E$ and $F$ are uniformly bounded away from zero. Since $E$ is uniformly contracted, by the classical stable manifold theory, for every $x\in\Lambda^{\prime}$ there is a well defined local stable manifold $W^{ss}_{loc}(x)$ with uniform size along $\Lambda^{\prime}$.
Since $\Lambda^{\prime}$ is not hyperbolic, there may not be well defined for local unstable manifolds for every point in $\Lambda^{\prime}$. Moreover, even though these local unstable manifolds may exist for some points (e.g. along non-singular compact and invariant subsets) they do not need to have uniform size. To overcome this issue, we firstly recall that the periodic orbits $\gamma_{n}$ are hyperbolic by the hyperbolic lemma. So, fix $\gamma_{n}$ and let
$T\gamma_{n}=E^{\prime s}\oplus\langle X\rangle\oplus E^{\prime u}$
be the hyperbolic splitting of $\gamma_{n}$. Since $E$ is uniformly contracted, we have $E(x)\subset E^{\prime s}(x)$, for every $x\in\gamma_{n}$. Now, suppose there is $x\in\gamma_{n}$ so that $E^{\prime s}(x)\not\subset E(x)$. In this case, there is a vector $v\in E^{\prime s}(x)$ such that $v=v^{E}+v^{F}$, $v^{E}\in E(x)$, $v^{F}\in F(x)$ and $v^{F}\neq 0$. Moreover, since $v^{F}\in E^{\prime s}(x)$, then $v\notin\langle X(x)\rangle$. On the other hand, by considering the subspace $L\subset F$ generated by $v^{F}$ and $\langle X(x)\rangle$, the sectional-expansion of $L$ implies, $v^{F}\notin E^{\prime s}(x)$. This contradiction implies that $E(x)=E^{\prime s}(x)$ for every $x\in\gamma_{n}$. Consequently,
$\langle X(x)\rangle\oplus E^{\prime u}(x)=F(x),\quad\forall x\in\gamma_{n}.$
Choose $r>0$ as in Lemma 3.8 such that $B_{r}(\sigma_{X})\cap Sing(X)=\{\sigma_{X}\}$. Since any singularity contained in $C(\sigma_{X})$ is hyperbolic by Lemma 2.2, there are only finitely many of them. By shrinking $U$ if it is necessary, we have that $Sing(X|_{U})=Sing(X|_{C(\sigma_{X})})$ and hence, there is a neighborhood $V$ of $Sing(X|_{C(\sigma_{X})})$ contained in $U$ satisfying
  • $V_{i}=B_{r}(\sigma_{j})$, where $\{\sigma_{j}:j=1,\ldots,k\}=Sing(X|_{C(\sigma_{X})})$.
  • $V=V_{1}\cup\cdots\cup V_{j}$.
  • Each $V_{i}$ is a neighborhood of a unique singularity in $X|_{U}$.
Fix $T_{0}>0$ in a such way that
(3) $Ke^{-\lambda t}\leq e^{-\lambda^{\prime}t}\text{ and }K_{0}^{2}e^{-\lambda^{\prime}t}\leq e^{-\lambda^{\prime\prime}t},\quad\forall t\geq T_{0},$
where $K,\lambda>0$ the constants given by sectional hyperbolicity of $\Lambda^{\prime}$, $0<\lambda^{\prime\prime}<\lambda^{\prime}<\lambda$ and
$K_{0}=\sup\left\{\frac{\|X(x)\|}{\|X(y)\|}:x,y\in C(\sigma_{X})\setminus V\right\}.$
For any point $q\in V$, let $q_{e}$ be given as (2). Suppose that $\phi_{[-T,0]}(q^{n}_{e})\subset C(\sigma_{X})\setminus V$, where $T>T_{0}$. Define $\eta_{0}=e^{-\lambda^{\prime\prime}T_{0}}$.
Claim 2.
The orbit segment $\phi_{[-T,0]}(q_{e})$ is $(\eta_{0},T_{0})$-backward contracting for the bundle $\mathcal{F}$.
Proof.
Let $0=t_{0}<t_{1}<\cdots<t_{m}=T$ such that $t_{j+1}-t_{j}=T_{0}$ for every $i=0,\ldots,m-2$ and $t_{m}-t_{m-1}\in(T_{0},2T_{0})$. Notice that the orbit segment $\phi_{[-T,0]}(q_{e})\subset C(\sigma_{X})\setminus V$. Then, for every unitary vector $v\in\mathcal{F}_{q_{e}}$ one has
$Ke^{-\lambda(t_{j+1}-t_{j})}$ $\geq$ $|\det\Phi_{-(t_{j+1}-t_{j})}(\phi_{t_{-j}}(q_{e}))|$
$\geq$ $\frac{\|X(X_{-t_{j}}(q_{e}))\|}{\|X(X_{-t_{j+1}}(q_{e}))\|}\cdot\|\psi_{-(t_{j+1}-t_{j})}(\phi_{t_{-j}}(q_{e}))v_{j}\|,$
where $v_{1}=v$ and $v_{j+1}=\psi_{-(t_{j+1}-t_{j})}(X_{-t_{j}}(q_{n}))v_{j}$ for any $j=1,\ldots,m-1$, and so, by (3),
$\|\psi_{-(t_{j+1}-t_{j})}(X_{-t_{j}}(q_{e}))|_{\mathcal{F}_{X_{-t_{j}}(q_{e})}}\|\leq\frac{\|X(X_{-t_{j+1}}(q_{e}))\|}{\|X(X_{-t_{j}}(q_{e}))\|}e^{-\lambda^{\prime}(t_{j+1}-t_{j})},\quad\forall j=0,\ldots,m-1.$
Thus, we have by the choice of $K_{0}$ in (3) that
$\prod_{j=0}^{m-1}\|\psi^{*}_{-(t_{j+1}-t_{i})}(X_{-t_{j}}(q_{e}))|_{\mathcal{F}_{X_{-t_{j}}(q_{e})}}\|$ $\leq$ $\prod_{j=0}^{m-1}\left[\frac{\|X(X_{-t_{j+1}}(q_{e}))\|}{\|X(X_{-t_{j}}(q_{e}))\|}\right]^{2}e^{-\lambda^{\prime}(t_{i+1}-t_{i})}$
$\leq$ $\left[\frac{\|X(X_{-t_{1}}(q_{e}))\|}{\|X(q_{e})\|}\right]^{2}e^{-m\lambda^{\prime}T_{0}}$
$\leq$ $\left(K_{0}\right)^{2}e^{-m\lambda^{\prime}T_{0}}$
$\leq$ $(e^{-\lambda^{\prime\prime}T_{0}})^{m}=\eta_{0}^{m}.$
This proves the claim. ∎
Notice that if $T_{0}>T_{1}>0$ and $0<\lambda_{1}<\lambda^{\prime\prime}$, then
$\lambda^{\prime\prime}T_{0}=\lambda^{\prime\prime}T_{1}\cdot\frac{T_{0}}{T_{1}}>\lambda^{\prime\prime}T_{1}$ $\Rightarrow$ $-\lambda^{\prime\prime}T_{0}<-\lambda^{\prime\prime}T_{1}<-\lambda_{1}T_{1}$
$\Rightarrow$ $\eta_{0}=e^{-\lambda^{\prime\prime}T_{0}}<e^{-\lambda^{\prime\prime}T_{1}}<e^{-\lambda_{1}T_{1}}=\eta^{\prime}.$
Hence, if the orbit segment $\phi_{[-T,0]}(q_{e})$ is $(\eta_{0},T_{0})$-backward contracting for the bundle $\mathcal{F}$, then it is also $(\eta_{1},T_{1})$-backward contracting for the same bundle $\mathcal{F}$ for every $0<T_{1}<T_{0}$ and $\eta\in(\eta^{\prime},1)$.
Since $\Pi(\gamma_{n})\to+\infty$, we have by te Pliss Lemma that there are $\eta,\eta_{1},T_{1}>0$ and $p_{n}\in\gamma_{n}$ such that $p_{n}$ is a $(\eta,T_{1})$-backward hyperbolic point for $\mathcal{F}$ along the orbit of $\gamma_{n}$. Moreover, such points have positive density $a>0$ along the orbit of $\gamma_{n}$ for every $n\geq 1$. In this case, we have the following possibilities:
  • $\sup\{T>0:X_{[-T,0]}(q_{e,n})\subset C(\sigma_{X})\setminus V\}<T_{0}$ for every $n\geq 1$. Since the density of $(\eta,T_{1})$-backward hyperbolic times is $a>0$ it follows that there must be hyperbolic times $p^{\prime}_{n}\in\gamma_{n}\cap V_{\sigma_{X}}$. In particular, there is $s_{n}\in\mathbb{R}$ such that $X_{s_{n}}(p_{n})\in\overline{D^{s}_{\alpha}(\sigma_{X})}$, or
  • there is a subsequence $\{p_{n_{k}}\}_{k\geq 1}$ such that
    $\sup\{T>0:X_{[-T,0]}(q_{e,n_{k}})\subset C(\sigma_{X})\setminus V\}\geq T_{0},\quad\forall k\geq 1.$
    In this case, if $p_{n_{k}}\in V_{\sigma_{X}}$, we proceed as item 1 above. If $p_{n_{k}}\notin V_{\sigma_{X}}$, then by the Claim 2 and (3), there is $s>0$ such that $X_{s}(p_{n_{k}})=q^{n_{k}}_{e}\in\overline{D^{s}_{\alpha}(\sigma)}$ is also a $(\eta,T_{1})$-backward hyperbolic point for the bundle $\mathcal{F}$ for every $k\geq 1$.
So, it follows from the both cases and by taking a sequence if it is necessary, that $q_{n}\in\gamma_{n}$ is a $(\eta,T_{1})$ backward hyperbolic points inside $\overline{D^{s}_{\alpha}(\sigma)}$ for any $n\geq 1$. Therefore, by Lemma 3.8 we have that
$W^{u}(p_{n})\cap W^{s}(\sigma)\neq\varnothing.$
@@
Let $\Lambda=C(\sigma_{X})$ and $\Lambda_{Y}=C(\sigma_{Y})$. Fix a neighborhood $U$ for $\Lambda$ and $\Lambda_{Y}$, given by the continuity of the map $Y\to C(\gamma_{Y})$. Since $\Lambda$ is partially hyperbolic, it is possible to extend the sectional-hyperbolic splitting $T_{\Lambda}M$ to $U$ (by shrinking $U$ if it is possible), that we will keep written as $T_{U}M=E\oplus F$. It should be noted that since $\Lambda$ is sectional-hyperbolic one has $1\leq\dim E\leq n-2$ and $1\leq\dim F\leq n-1$. Thus, the subspace $N_{x}$ is of codimension one and has a coordinate system induced by the sectional-hyperbolic splitting for every $x\in U$ in the following way:
$\mathcal{E}_{x}=\pi(E_{x})\text{ and }\mathcal{F}_{x}=\pi(F)\text{ or }F_{x}\cap N_{x},\quad x\in U,$
From partial hyperbolicity we have that the angle between these subspaces is uniformly bounded away from zero. Denote by $N(\eta)$ the standard cube $(-\eta,\eta)^{n-1}\subset N_{x}$, according to this coordinate system.
For any regular point $x\in U$, there is a number $\eta(x)>0$ such that the exponential map $Exp_{x}:N_{x}\to M$ is an isometry. Note that for every $\delta>0$, there is $\eta_{\delta}>0$ such that $\eta(x)>\eta_{\delta}$ for any $x\in U\setminus B(Sing(\Lambda),\delta)$. Let $\eta^{*}:U\to\mathbb{R}^{+}$ defined by $\eta^{*}(x)=\min\{\eta(x),\eta_{\delta}\}$ for every $x\in U$. Denote by $\mathcal{N}_{x}=Exp_{x}(N_{x}(\eta^{*}(x)))\subset M$, a codimension 1 cross section on $x$ of size $\eta^{*}(x)$.
For a regular point $x$, the coordinates $\mathcal{E}_{x}$ and $\mathcal{F}_{x}$ allows us to define the volume of boxes $B\subset N_{x}$ with respect to these axis as follows: Denote by $\hat{\pi}_{s}:N_{x}\to\mathcal{E}_{x}$ and $\hat{\pi}_{c}:N_{x}\to\mathcal{F}_{x}$ the projections along the $\mathcal{E}_{x}$-axis and the $\mathcal{F}_{x}$-axis respectively. The stable (central) volume of $B$ is given by
@@
where $vol_{s(c)}(A)$ denotes the volume of a subset $A\subset\mathbb{R}^{\dim\mathcal{E}_{x}(\mathcal{F}_{x})}$. We shall use a vector $v=(\varepsilon^{s}_{1},\varepsilon^{s}_{2},\ldots,\varepsilon^{s}_{n_{s}},\varepsilon^{c}_{1},\varepsilon^{s}_{2},\ldots,\varepsilon^{c}_{n_{c}})$, where $1\leq n_{s}\leq n-2$ and $1<n_{c}\leq n-1$ satisfy $n_{s}+n_{c}+1=n$, to describe the box $B[v]$ with stable volume $\varepsilon_{s}$ and central volume $\varepsilon_{c}$.
@@
$H_{x}^{t}:dom(H_{x}^{t})\subset\mathcal{N}_{x}\to\mathcal{N}_{X_{t}(x)},$